Problem 9-1A Short-term notes payable transactions and entries LO P1 [The following information applies to the questions disp[The following information applies to the questions displayed below.) Tyrell Co. entered into the following transactions invoProblem 9-1A Short-term notes payable transactions and entries LO P1 [The following information applies to the questions dispProblem 9-1A Short-term notes payable transactions and entries LO P1 [The following information applies to the questions dispProblem 9-1A Part 5 5. Prepare journal entries for all the preceding transactions and events. (Do not round your intermediate
Answer #1
1) Maturity date
locust NBR fargo
date of the note 19-May 8-Jul 28-Nov
term of note 90 120 60
maturity date 17-Aug 5-Nov 27-Jan
2) interest due at maturity
principal * Rate * time = interest
locust 35,000 * 8% * 90/360 = 700
NBR 63,000 * 11% * 120/360 = 2310
Fargo 30,000 * 7% * 60/360 = 350
3) Amount in adjusting entry
Fargo Bank
principal * Rate * time = interest
interest to be acccrued in 2016 30,000 * 7% * 33/360 = 193
4) interest expense to be recorded in 2017
principal * Rate * time = interest
interest to recorded in 2018 30,000 * 7% * 27/360 = 158
Journal entries
Date Accounting titles & Explanations Debit Credit
2016
20-Apr inventory 35,500
Accounts payable 35,500
19-May Accounts payable 35,500
cash 500
notes payable 35,000
8-Jul Cash 63,000
notes payable 63,000
17-Aug notes payable 35,000
interest expense 700
cash 35,700
5-Nov notes payable 63,000
interest expense 2,310
cash 65,310
28-Nov Cash 30,000
notes payable 30,000
31-Dec interest expense 193
interest payable 193
2017
27-Jan notes payable 30,000
interest payable 193
interest expense 158
cash 30,350